### 06 Apr 2026 v1

@@ 5 Complexity analysis
Let $h_{x}=\mathcal{B}h_{\hat{x}}$ and $\hat{h}=(\hat{h}_{\hat{x}},\hat{h}_{s})=(\mathcal{H}^{-1}h_{x},\mathcal{H}^{-1}h_{s})$. Define
$\omega(\mu):=S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h].$
@@
where
$\omega_{x}(\mu):=\rho\langle h_{x},(\mathcal{I}_{\mathbb{E}}-\mathcal{H}^{-1})h_{x}\rangle,\quad\omega_{s}(\mu):=\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle.$
Differentiating $\omega_{x}(\mu)$ with respect to $\mu$ yields
$\omega^{\prime}_{x}(\mu)$ $=\langle\mathcal{H}^{-1}h_{x},D^{2}\phi(z_{\rho})\mathcal{H}^{-1}h_{x}\rangle+\mu D^{3}\phi(z_{\rho})[z_{\rho}^{\prime},\mathcal{H}^{-1}h_{x},\mathcal{H}^{-1}h_{x}]$ (114)
$=\frac{\rho}{\mu}\langle h_{x},\mathcal{H}^{-1}\mathcal{W}\mathcal{H}^{-1}h_{x}\rangle+\mu D^{3}\phi(z_{\rho})[z_{\rho}^{\prime},\hat{h}_{\hat{x}},\hat{h}_{\hat{x}}$]
$\overset{(i)}{\leq}\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+2\mu\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}},\hat{h}_{\hat{x}}$]
$=\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+2\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}\left<\mathcal{H}^{-1}h_{x},\mathcal{W}\mathcal{H}^{-1}h_{x}\right$>
$\overset{(ii)}{\leq}\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+2\sqrt{\frac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$]
$\overset{(iii)}{\leq}\frac{\rho}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]+\frac{2\rho}{\mu}\sqrt{\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$]
$\overset{(iv)}{=}\dfrac{\rho(1+2\sqrt{\nu})}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$].
Here, the inequalities (i)(i)$(ii)$ follow from Corollary 1, Theorem 3.1, the 11-self-concordance of $\phi$, and $\mathcal{H}^{-1}\prec\mathcal{I}_{\mathbb{E}}$. The inequality $(iii)$ uses $\mathcal{H}^{-1}\prec\mathcal{W}^{-1}$. The final identity $(iv)$ is due to (11).
@@
$\omega^{\prime}_{s}(\mu)\leq-\frac{1+2\sqrt{\nu}}{\mu}D^{2}_{ss}\eta_{\rho}[h_{s},h_{s}].$ (115)
Combing (114) and (115) gives
$\omega^{\prime}(\mu)$ $\leq\dfrac{\rho(1+2\sqrt{\nu})}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]-\frac{1+2\sqrt{\nu}}{\mu}D^{2}_{ss}\eta_{\rho}[h_{s},h_{s}]$ (116)
$\leq\frac{\rho(1+{2\sqrt{\nu}})}{\mu}S_{\eta_{\rho}}[h,h],$
which completes the proof.
The following theorem provides an upper bound for the primal-dual gap function $\theta_{\rho}(\hat{x},s;\mu)$ that depends only on $\kappa$ and $\mu$.
@@
For any $\mu>0$, $\rho\geq 1$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$, suppose that
$\xi_{\rho}(\hat{x},s;\mu)\leq\kappa.$
@@
To quantify the gaps between the current value $\eta_{\rho}$ and the primal and dual optimal values, define
$\theta_{\hat{x},\rho}(\mu)$ $=\eta_{\rho}(\hat{x},{s};\mu)-\eta_{\rho}(\hat{x}_{\rho}(s,\mu),{s};\mu),$ (118)
@@
If $\xi_{\rho}(\hat{x},s;\mu)\leq\kappa$, then by Theorem 2.3(v),
$\max\left\{\tilde{\delta}_{\hat{x},\rho}(\hat{x},s;\mu),\tilde{\delta}_{s,\rho}(\hat{x},s;\mu)\right\}\leq 2\kappa.$ (119)
@@
$\theta_{s,\rho}(\mu)$ $=\eta_{\rho}(\hat{x},{s}(\hat{x},\mu);\mu)-\eta_{\rho}(\hat{x},{s};\mu)$ (120)
$=-\int_{0}^{1}\langle\widetilde{\Delta s},\nabla_{s}\eta_{\rho}(\hat{x},{s}(\hat{x},\mu)+\tau\widetilde{\Delta s};\mu)\rangle\;d\tau$
$=-\int_{0}^{1}\int_{0}^{\tau}\langle\widetilde{\Delta s},D^{2}_{ss}\eta_{\rho}(\hat{x},{s}(\hat{x},\mu)+t\widetilde{\Delta s};\mu)\widetilde{\Delta s}\rangle\;dtd\tau$
@@
$\theta_{\rho}(\hat{x},s;\mu)=\theta_{s,\rho}(\mu)+\theta_{\hat{x},\rho}(\mu)\leq\kappa\mu.$ (121)
@@
By (76) and Theorem 5.4, if $\xi_{\rho}(\hat{x},s;\mu)\leq\kappa$, then
$|\eta_{\rho}(\hat{x},s;\mu)-{\rm val}(\mathrm{P}_{\mu})|\leq\kappa\mu.$
Therefore, the PFSNM can be viewed as a relaxation of the IPM. The interpolation between the subproblem objectives of the two methods is controlled by the parameter $\mu$, which establishes a direct connection between the SNM and the IPM.
The following lemma relates the merit function $\xi_{\rho}(\hat{x},s;\mu)$ to the quantities $\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)$ and $S_{\eta_{\rho}}(\hat{x},s;\mu)$, which is crucial for the subsequent complexity analysis.
@@
For any $\mu>0$, $\rho\geq 1$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\xi_{\rho}(\hat{x},s;\mu)=\max\limits_{0\neq h\in\hat{\mathbb{E}}\times\mathbb{E}}\frac{\left|\langle\nabla_{w}\eta_{\rho}(\hat{x},s;\mu),h\rangle\right|}{\sqrt{\mu S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]}}.$ (122)
@@
$\sqrt{\mu}\xi_{\hat{x},\rho}(\hat{x},s;\mu)\geq\frac{\left|\langle\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu),h_{\hat{x}}\rangle\right|}{\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)h_{\hat{x}}\rangle}}$, (123)
$\sqrt{\mu}\xi_{s,\rho}(\hat{x},s;\mu)\geq\frac{\left|\langle\nabla_{s}\eta_{\rho}(\hat{x},s;\mu),h_{s}\rangle\right|}{\sqrt{-\langle h_{s},D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)h_{s}\rangle}}.$
@@
$\sqrt{\mu}\xi_{\rho}\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}h_{\hat{x}}\rangle-\langle h_{s},D^{2}_{ss}\eta_{\rho}h_{s}\rangle}$ (124)
@@
$\overset{(i)}{\geq}$ $\ \sqrt{\mu}\left(\xi_{\hat{x},\rho}\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}h_{\hat{x}}\rangle}+\xi_{s,\rho}\sqrt{-\langle h_{s},D^{2}_{ss}\eta_{\rho}h_{s}\rangle}\right)$
$\overset{(ii)}{\geq}$ $\left|\langle\nabla_{\hat{x}}\eta_{\rho},h_{\hat{x}}\rangle\right|+\left|\langle\nabla_{s}\eta_{\rho},h_{s}\rangle\right|$
@@
where both $(i)$ and $(ii)$ follow from the Cauchy–Schwarz inequality. Thus,
$\xi_{\rho}(\hat{x},s;\mu)\geq\frac{|\langle\nabla_{w}\eta_{\rho}(\hat{x},s;\mu),h\rangle|}{\sqrt{\mu S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]}},\quad\forall\,h\in\hat{\mathbb{E}}\times\mathbb{E},h\neq 0.$ (125)
@@
$h=\left((D_{\hat{x}\hat{x}}^{2}\eta_{\rho})^{-1}\nabla_{\hat{x}}\eta_{\rho},(D_{ss}^{2}\eta_{\rho})^{-1}\nabla_{s}\eta_{\rho}\right),$ (126)
@@
Building on these estimates, we establish the main result of the paper: Algorithm 2 admits a polynomial-time complexity bound of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the best‐known complexity of the classical short-step IPMs.
@@
Let $\rho\geq 1$ and choose
$1>\sigma\geq 1-\dfrac{\ln(\gamma)}{2\rho\sqrt{\nu}+\ln(\gamma)},\;where\;\gamma:=\dfrac{2\kappa+\sqrt{2}/\rho}{\kappa+\sqrt{2}/\rho}.$ (127)
Then Algorithm 2 requires at most $\mathcal{O}(\sqrt{\nu}\ln(\mu^{(0)}/\varepsilon))$ iterations to attain the desired accuracy $\varepsilon$.

### 21 May 2026 v2

@@ 5 Complexity analysis
Let $h_{x}=\mathcal{B}h_{\hat{x}}$ and $\bar{h}=(\bar{h}_{x},\bar{h}_{s})=(\mathcal{H}^{-1}h_{x},\mathcal{H}^{-1}h_{s})$. Define
$\omega(\mu):=S_{\eta_{\rho}}(w;\mu)[h,h].$
@@
where $\omega_{x}(\mu):=\rho\langle h_{x},(\mathcal{I}_{\mathbb{E}}-\mathcal{H}^{-1})h_{x}\rangle,\,\omega_{s}(\mu):=\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle.$ Differentiating $\omega_{x}(\mu)$ with respect to $\mu$ yields
$\left|\omega^{\prime}_{x}(\mu)\right|$ $=\left|\langle\mathcal{H}^{-1}h_{x},D^{2}\phi(z_{\rho})\mathcal{H}^{-1}h_{x}\rangle+\mu D^{3}\phi(z_{\rho})[z_{\rho}^{\prime},\mathcal{H}^{-1}h_{x},\mathcal{H}^{-1}h_{x}]\right|$ (78)
$\leq\left|\frac{\rho}{\mu}\langle h_{x},\mathcal{H}^{-1}\mathcal{W}\mathcal{H}^{-1}h_{x}\rangle\right|+\left|\mu D^{3}\phi(z_{\rho})[z_{\rho}^{\prime},\bar{h}_{x},\bar{h}_{x}]\right|.$
By Corollary 1 and the inequality $\mathcal{H}^{-1}\mathcal{W}\mathcal{H}^{-1}\prec\mathcal{H}^{-1}\mathcal{W}$, we have
$\left|\rho\langle h_{x},\mathcal{H}^{-1}\mathcal{W}\mathcal{H}^{-1}h_{x}\rangle\right|\leq D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$]. (79)
Since $\phi$ is a standard self-concordant convex function, it follows from (Nesterov and Nemirovskii, 1994, Proposition 9.1.1) and (79) that
$\left|\mu D^{3}\phi(z_{\rho})[z_{\rho}^{\prime},\bar{h}_{x},\bar{h}_{x}]\right|$ $\leq 2\mu\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}D^{2}\phi(z_{\rho})[\bar{h}_{x},\bar{h}_{x}]$ (80)
$=2\rho\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}\left<h_{x},\mathcal{H}^{-1}\mathcal{W}\mathcal{H}^{-1}h_{x}\right$>
$\leq 2\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$].
By Theorem 3.1 and the inequality $\mathcal{H}^{-1}\prec\mathcal{W}^{-1}$, we have
$\sqrt{D^{2}\phi(z_{\rho})[z_{\rho}^{\prime},z_{\rho}^{\prime}]}$ $=\frac{1}{\rho}\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{H}^{-1}D^{2}\phi(z_{\rho})\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}$ (81)
$\leq\frac{1}{\rho}\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{W}^{-1}D^{2}\phi(z_{\rho})\mathcal{W}^{-1}\nabla\phi(z_{\rho})\rangle}$
$=\frac{1}{\mu}\sqrt{\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}$
$=\frac{\sqrt{\nu}}{\mu}.$
Combining (78)–(81) yields
$\left|\omega^{\prime}_{x}(\mu)\right|\leq\frac{1+2\sqrt{\nu}}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}$]. (82)
@@
$\left|\omega^{\prime}_{s}(\mu)\right|\leq\frac{1+2\sqrt{\nu}}{\mu}\left(-D^{2}_{ss}\eta_{\rho}[h_{s},h_{s}]\right).$ (83)
Since $\omega(\mu)=\omega_{x}(\mu)+\omega_{s}(\mu)$, it follows from the triangle inequality, (82), and (83) that
$|\omega^{\prime}(\mu)|$ $\leq\dfrac{1+2\sqrt{\nu}}{\mu}D^{2}_{\hat{x}\hat{x}}\eta_{\rho}[h_{\hat{x}},h_{\hat{x}}]-\frac{1+2\sqrt{\nu}}{\mu}D^{2}_{ss}\eta_{\rho}[h_{s},h_{s}]$
$=\frac{1+{2\sqrt{\nu}}}{\mu}S_{\eta_{\rho}}[h,h].$
This completes the proof.
The preceding two theorems quantify how the first- and second-order quantities induced by the reduced BAL function vary when the smoothing parameter is changed. These estimates will be used to control the effect of the parameter $\mu$ update in the path-following phase. Before deriving the final complexity bound, we also need an estimate that connects the computable quantity $\xi_{\rho}(w;\mu)$ with the primal-dual gap $\theta_{\rho}(w;\mu)$. For this purpose, we introduce auxiliary displacement measures from the current point to the exact partial minimizer and maximizer of the reduced minimax problem:
$\tilde{\delta}_{\hat{x},\rho}(w;\mu)$ $=\sqrt{\dfrac{1}{\mu}\langle\widetilde{\Delta\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(w;\mu)\widetilde{\Delta\hat{x}}\rangle}$, (84a)
$\tilde{\delta}_{s,\rho}(w;\mu)$ $=\sqrt{-\dfrac{1}{\mu}\langle\widetilde{\Delta s},D^{2}_{ss}\eta_{\rho}(w;\mu)\widetilde{\Delta s}\rangle}$, (84b)
$\tilde{\delta_{\rho}}(w;\mu)$ $=\sqrt{(\tilde{\delta}_{\hat{x},\rho}(w;\mu))^{2}+(\tilde{\delta}_{s,\rho}(w;\mu))^{2}}=\|\widetilde{\Delta w}\|_{\eta_{\rho},w},$ (84c)
[…]
@@
For any $\mu>0$, $\rho>0$, and any point $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$, suppose that $\kappa=0.1$ and
$\xi_{\rho}(w;\mu)\leq\kappa.$
@@
To quantify the gaps between the current value $\eta_{\rho}$ and the primal optimal value, define
$\theta_{\hat{x},\rho}(\mu)$ $=\eta_{\rho}(\hat{x},{s};\mu)-\eta_{\rho}(\hat{x}_{\rho}(s,\mu),{s};\mu),$
@@
If $\xi_{\rho}(w;\mu)\leq\kappa$, then by Theorem 2.3(iv),
$\max\left\{\tilde{\delta}_{\hat{x},\rho}(w;\mu),\tilde{\delta}_{s,\rho}(w;\mu)\right\}\leq 2\kappa.$
@@
$\theta_{s,\rho}(\mu)$ $=\eta_{\rho}(\hat{x},{s}(\hat{x},\mu);\mu)-\eta_{\rho}(\hat{x},{s};\mu)$
$=-\int_{0}^{1}\langle\widetilde{\Delta s},\nabla_{s}\eta_{\rho}(\hat{x},s_{\rho}(\hat{x},\mu)+\tau\widetilde{\Delta s};\mu)\rangle\;d\tau$
$=-\int_{0}^{1}\int_{0}^{\tau}\langle\widetilde{\Delta s},D^{2}_{ss}\eta_{\rho}(\hat{x},s_{\rho}(\hat{x},\mu)+t\widetilde{\Delta s};\mu)\widetilde{\Delta s}\rangle\;dtd\tau$
@@
$\theta_{\rho}(w;\mu)=\theta_{s,\rho}(\mu)+\theta_{\hat{x},\rho}(\mu)\leq\kappa\mu.$
@@
By (54) and Theorem 5.4, if $\xi_{\rho}(w;\mu)\leq\kappa$, then
$|\eta_{\rho}(w;\mu)-{\rm val}(\mathrm{P}_{\mu})|\leq\kappa\mu.$
Therefore, the PFSNM can be viewed as a relaxation of the IPM. The interpolation between the subproblem objectives of the two methods is controlled by the parameter $\mu$.
The following lemma relates the merit function $\xi_{\rho}(w;\mu)$ to the quantities $\nabla_{w}\eta_{\rho}(w;\mu)$ and $S_{\eta_{\rho}}(w;\mu)$, which is crucial for the subsequent complexity analysis.
@@
For any $\mu>0$, $\rho>0$, and any point $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\xi_{\rho}(w;\mu)=\max\limits_{0\neq h\in\hat{\mathbb{E}}\times\mathbb{E}}\frac{\left|\langle\nabla_{w}\eta_{\rho}(w;\mu),h\rangle\right|}{\sqrt{\mu S_{\eta_{\rho}}(w;\mu)[h,h]}}.$ (85)
@@
$\sqrt{\mu}\xi_{\hat{x},\rho}(w;\mu)\geq\frac{\left|\langle\nabla_{\hat{x}}\eta_{\rho}(w;\mu),h_{\hat{x}}\rangle\right|}{\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(w;\mu)h_{\hat{x}}\rangle}}$,
$\sqrt{\mu}\xi_{s,\rho}(w;\mu)\geq\frac{\left|\langle\nabla_{s}\eta_{\rho}(w;\mu),h_{s}\rangle\right|}{\sqrt{-\langle h_{s},D^{2}_{ss}\eta_{\rho}(w;\mu)h_{s}\rangle}}.$
@@
$\sqrt{\mu}\xi_{\rho}\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}h_{\hat{x}}\rangle-\langle h_{s},D^{2}_{ss}\eta_{\rho}h_{s}\rangle}$
@@
$\geq$ $\ \sqrt{\mu}\left(\xi_{\hat{x},\rho}\sqrt{\langle h_{\hat{x}},D^{2}_{\hat{x}\hat{x}}\eta_{\rho}h_{\hat{x}}\rangle}+\xi_{s,\rho}\sqrt{-\langle h_{s},D^{2}_{ss}\eta_{\rho}h_{s}\rangle}\right)$
$\geq$ $\ \left|\langle\nabla_{\hat{x}}\eta_{\rho},h_{\hat{x}}\rangle\right|+\left|\langle\nabla_{s}\eta_{\rho},h_{s}\rangle\right|$
@@
where the first inequality follows from the Cauchy–Schwarz inequality. Thus,
$\xi_{\rho}(w;\mu)\geq\frac{|\langle\nabla_{w}\eta_{\rho}(w;\mu),h\rangle|}{\sqrt{\mu S_{\eta_{\rho}}(w;\mu)[h,h]}},\quad\forall\,h\in\hat{\mathbb{E}}\times\mathbb{E},\,h\neq 0.$ (86)
@@
$h=\left((D_{\hat{x}\hat{x}}^{2}\eta_{\rho})^{-1}\nabla_{\hat{x}}\eta_{\rho},-(D_{ss}^{2}\eta_{\rho})^{-1}\nabla_{s}\eta_{\rho}\right),$
@@
The preceding estimates quantify the sensitivity of the reduced BAL function with respect to the smoothing parameter. The following lemma is the key step in the path-following analysis: it shows that, if the current point is in the neighborhood associated with $\mu$, then the same point remains sufficiently close to the new central path after replacing $\mu$ by $\sigma\mu$.
Lemma 4
Let $\rho>0$, $\mu>0$, $\kappa=0.1$, and $\mu^{+}=\sigma\mu$ with
$\sigma=1-\dfrac{\ln(\gamma)}{2\sqrt{\nu}+\ln(\gamma)},\;where\;\gamma:=\dfrac{2\kappa+\sqrt{2}}{\kappa+\sqrt{2}}.$ (87)
If $\xi_{\rho}(w;\mu)\leq\kappa$, then
$\xi_{\rho}(w;\mu^{+})\leq 2\kappa.$ (88)
Proof
For any nonzero $h\in\hat{\mathbb{E}}\times\mathbb{E}$, define the function $\psi:\mathbb{R}_{++}\times\hat{\mathbb{E}}\times\mathbb{E}\to\mathbb{R}$ as
$\psi(\tau,h):=\frac{\langle\nabla_{w}\eta_{\rho}(w;\tau),h\rangle^{2}}{\tau S_{\eta_{\rho}}(w;\tau)[h,h]}.$
Differentiating $\psi(\tau,h)$ with respect to $\tau$ gives
$\psi^{\prime}(\tau,h)$ $=\frac{2\langle\nabla_{w}\eta_{\rho}(w;\tau),h\rangle\cdot\langle\nabla_{w}\eta_{\rho}^{\prime}(w;\tau),h\rangle}{\tau S_{\eta_{\rho}}(w;\tau)[h,h]}-\frac{\langle\nabla_{w}\eta_{\rho}(w;\tau),h\rangle^{2}}{\tau^{2}S_{\eta_{\rho}}(w;\tau)[h,h]}$
$\qquad\qquad-\frac{\langle\nabla_{w}\eta_{\rho}(w;\tau),h\rangle^{2}S^{\prime}_{\eta_{\rho}}(w;\tau)[h,h]}{\tau(S_{\eta_{\rho}}(w;\tau)[h,h])^{2}}.$
Let $\tau\in[\mu^{+},\mu]$. It follows from Theorems 5.25.3 that
$|\psi^{\prime}(\tau,h)|$ $\leq\sqrt{\frac{8\nu}{\tau^{2}}}\sqrt{\psi(\tau,h)}+\frac{2+2\sqrt{\nu}}{\tau}\psi(\tau,h)$
$\leq\sqrt{\frac{8\nu}{(\mu^{+})^{2}}}\sqrt{\psi(\tau,h)}+\frac{2+2\sqrt{\nu}}{\mu^{+}}\psi(\tau,h).$
Define $c_{1}:=\frac{1+\sqrt{\nu}}{\mu^{+}}$ and $\Psi(\tau,h):=e^{c_{1}\tau}\sqrt{\psi(\tau,h)}$. We have
$\Psi^{\prime}(\tau,h)\leq e^{c_{1}\tau}\frac{\sqrt{2\nu}}{\mu^{+}}.$ (89)
Integrating both sides of (89) over $\mu^{+},\mu]$ yields
$\Psi(\mu^{+},h)-\Psi(\mu,h)\leq\frac{\sqrt{2\nu}}{\mu^{+}}\frac{1}{c_{1}}\left(e^{c_{1}\mu}-e^{c_{1}\mu^{+}}\right).$
$\sqrt{\psi(\mu^{+},h)}$ (90)
$\leq$ $\,e^{c_{1}(\mu-\mu^{+})}\sqrt{\psi(\mu,h)}+\frac{\sqrt{2\nu}}{\mu^{+}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu-\mu^{+})}-1\right)$
$\leq$ $\,e^{c_{1}(\mu-\mu^{+})}\xi_{\rho}(w;\mu)+\frac{\sqrt{2\nu}}{\mu^{+}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu-\mu^{+})}-1\right),$
$\xi_{\rho}(w;\mu^{+})$ $\leq e^{c_{1}(\mu-\mu^{+})}\xi_{\rho}(w;\mu)+\frac{\sqrt{2\nu}}{\mu^{+}}\frac{1}{c_{1}}\left(e^{c_{1}(\mu-\mu^{+})}-1\right).$ (91)
Recall $\mu^{+}=\sigma\mu$. Let
$\left\{\begin{array}[]{ll}&\alpha_{1}(\sigma):=e^{c_{1}(\mu-\mu^{+})}=e^{(1+\sqrt{\nu})(\frac{1}{\sigma}-1)},\\ &\alpha_{2}:=\frac{\sqrt{2\nu}}{\mu^{+}}\frac{1}{c_{1}}=\frac{\sqrt{2\nu}}{1+\sqrt{\nu}}.\end{array}\right$. (92)
It can be verified that
[…]
whenever $\sigma=1-\dfrac{\ln(\gamma)}{2\sqrt{\nu}+\ln(\gamma)}$.
Combining this estimate with (91) and (92), we get
$\xi_{\rho}(w;\mu^{+})\leq\alpha_{1}(\sigma)\kappa+\alpha_{2}(\alpha_{1}(\sigma)-1)\leq 2\kappa,$
which completes the proof.
Building on these estimates, we establish the main result of the paper: Algorithm 2 admits a polynomial iteration complexity of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the best–known complexity of classical short-step IPMs.
@@
Let $\rho>0$, $\kappa=0.1$, and choose $\sigma$ as defined in (87). Suppose that the initial point $v^{(0)}$ is generated by Algorithm 1. Then Algorithm 2 requires at most $\mathcal{O}(\sqrt{\nu}\ln(\mu^{(0)}/\varepsilon))$ iterations to attain the desired accuracy $\varepsilon$.